# TPV: Parameter Perturbations Through the Lens of Test Prediction Variance

### v2

@@ 2. Vary Number of Samples Experiment Details:
F.2 Details of Experiments in Section 5.2
F.2.1 Synthetic Data Experimental setup
@@
F.2.2 CIFAR experimental setup
@@
Noise-trained model. Each replicate begins by re-initializing the model to the clean reference weights $w^{\star}$. We then fine-tune this model for $10$ epochs using mini-batch SGD MSE regression on the noisy logits with momentum $0.9$, learning rate $10^{-4}$, batch size $256$, and 0 weight decay. Mini-batch shuffling is turned off and a random seed is used so that each run sees the same sequence of mini-batches in order to avoid randomness due to SGD and focus only on randomness due to label noise. Also, we train the models in Pytorch eval mode so that modules like batch norm do not make output logits batch dependent. See Appendix C.3.4 for more details.
@@
F.2.3 ImageNet experimental setup
For ImageNet, we use pre-trained resnet50, resnet18, wide_resnet50_2, shufflenet_v2_x1_0, efficientnet_b0, mnasnet1_0, convnext_tiny from Pytorch Hub. We randomly subsample $n_{\text{train}}=10000$ and $n_{\text{test}}=10000$. We use Gaussian noise level $\sigma=0.01$, $R=5$, and fine-tune this model for $10$ epochs using mini-batch SGD MSE regression on the noisy logits with momentum $0.9$, learning rate $10^{-6}$, batch size $64$, and 0 weight decay. While training, the loss is computed by summing the squared error over all the samples and logits in the mini-batch, but we track the MSE over samples. Additionally, it was extremely difficult to make the training loss go lower than the initial loss, and we had to turn on mini-batch shuffling within each run to make the loss go down. Nonetheless, we use a fixed random seed across runs so that the shuffling order remained the same across runs. We also tried experiments with pre-trained models mobilenet_v2, densenet121, and squeezenet1_0, but we were unable to make the MSE training loss go down for these models. All other details are identical to the above CIFAR experiments.
F.3 Pruning Experiment Details
@@
This appendix provides the experimental details for the pruning results reported in Section 5.3. All pruning experiments follow the global structured pruning protocol introduced in Optimal Brain Compression (OBC) (Chen et al., 2025). Our implementation is based on the official GitHub code of (Chen et al., 2025) and extends the OBC framework by adding the TPV-motivated JBR importance criteria.
F.3.1 Pruning Framework
@@ F.3.2 Models and Datasets
F.3.3 Pruning Criteria Compared
@@
F.3.4 Iterative Pruning Schedule

### v3

@@ 2. Vary Number of Samples Experiment Details:
G.2 Details of Experiments in Section 5.2
G.2.1 Synthetic Data Experimental setup
@@
G.2.2 CIFAR experimental setup
@@
Noise-trained model. Each replicate begins by re-initializing the model to the clean reference weights $w^{\star}$. We then fine-tune this model for $10$ epochs using mini-batch SGD MSE regression on the noisy logits with momentum $0.9$, learning rate $10^{-4}$, batch size $256$, and 0 weight decay. Mini-batch shuffling is turned off and a random seed is used so that each run sees the same sequence of mini-batches in order to avoid randomness due to SGD and focus only on randomness due to label noise. Also, we train the models in Pytorch eval mode so that modules like batch norm do not make output logits batch dependent. See Appendix D.3.4 for more details.
@@
G.2.3 ImageNet experimental setup
For ImageNet, we use pre-trained resnet50, resnet18, wide_resnet50_2, shufflenet_v2_x1_0, efficientnet_b0, mnasnet1_0, convnext_tiny from Pytorch Hub. We randomly subsample $n_{\text{train}}=10000$ and $n_{\text{test}}=10000$. We use Gaussian noise level $\sigma=0.01$, $R=5$, and fine-tune this model for $10$ epochs using mini-batch SGD MSE regression on the noisy logits with momentum $0.9$, learning rate $10^{-6}$, batch size $64$, and 0 weight decay. While training, the loss is computed by summing the squared error over all the samples and logits in the mini-batch, but we track the MSE over samples. Additionally, it was extremely difficult to make the training loss go lower than the initial loss, and we had to turn on mini-batch shuffling within each run to make the loss go down. Nonetheless, we use a fixed random seed across runs so that the shuffling order remained the same across runs. Thus the only source of stochasticity across runs was only the label noise. We also tried experiments with pre-trained models mobilenet_v2, densenet121, and squeezenet1_0, but we were unable to make the MSE training loss go down for these models. All other details are identical to the above CIFAR experiments.
Appendix H Pruning
H.1 Jacobian-Based Rebalancing (JBR) for Pruning
$u(x;w):=-\log(p_{c(x)}(x;w)),$
Pruning as a Source of Parameter Noise
$\delta w_{g}=\sigma\,\xi_{g}\,w_{g},\qquad g\in\mathcal{G},$ (75)
Connection and Contrast with JC
H.2 Related Work on Pruning
H.3 Pruning Experiments
We evaluate whether the TPV-motivated pruning criterion (JBR) improves accuracy–compression tradeoffs relative to standard groupwise criteria. Following the OBC pruning protocol, we perform global channel pruning with iterative removal and recomputation of importance scores. We compare JBR against Jacobian, L1-norm, BatchNorm-scale, FPGM, WHC, Taylor, and Random. We prune two ImageNet models (ResNet-50, MobileNet-V2) at $50\%$ global sparsity and two CIFAR models (ResNet-56 on CIFAR-10, VGG-19-BN on CIFAR-100) at $90\%$ sparsity. Each model is pruned iteratively for 18 steps, averaging results over 5 independent runs. Results are shown in Fig. 13 and Fig. 14 in terms of MACs. Across all architectures, the JBR criterion consistently matches or exceeds the performance of all baselines. See Appendix H.4 for details.
@@ H.3 Pruning Experiments
H.4 Pruning Experiment Details
This appendix provides the experimental details for the pruning results reported in Section H.3. All pruning experiments follow the global structured pruning protocol introduced in Optimal Brain Compression (OBC) (Chen et al., 2025). Our implementation is based on the official GitHub code of (Chen et al., 2025) and extends the OBC framework by adding the TPV-motivated JBR importance criteria.
H.4.1 Pruning Framework
@@ H.4.2 Models and Datasets
H.4.3 Pruning Criteria Compared
@@
H.4.4 Iterative Pruning Schedule